Project

General

Profile

February 21st

Date:
02/21/2017

Attendees:
Anna M., joe B., Vladimir P., Steve W., Margherita V.W., Marc M., D., Felipe A.

Agenda:
  1. POMS v2_0_0 status (All)
  2. Opened tickets (All)
  3. Alex’s requests (Felipe)
  4. AOB

Discussion:

1. POMS v2_0_0 status

The plan is to have the release next Wednesday, March 1st.

We all agree to move the testing task #15058 and all related sub-tasks to a later
release, v2.1.0 possibly 2 weeks after v2.0.0 is deployed.

FELIPE:

Task #15345 and Task #15317 are subtasks of #15058, move to v2.1.0

Task #15160: it's in draft mode but will be finished.

Task #15163: Almost done but will be moved to 2.1.0.
The plan is add also git flow procedure. Joe added the SSO authentication to the documentation for developers.
Vladimir also removed the use of the passw.ini file using pgpass.

Task #15529: Need to find a general way so we don't need to hardcode general definitions.
We have the poms.ini file.
Anna: the main reason for this request is to generalize the work environment in a unique way.

There was a long discussion about building , packaging and deploying the code: among others: makefile (Steve), software collection (Marc), etc.. The idea is to get away from ups/upd installation of the products

JOE:

Task #14248 It will be finished for the release.

Task #14377 Move to a future release. Still needs to understand discrepancy between elastic search and POMS in some numbers and see if we are losing data. We need to investigate more.

MARC:

Task #14274 Basically done. Need to update config of web server in production, but it will be
done at the prd release time.

Task #15058 Move this “testing task” to v2.1.0

Task #15332 Mostly done but it appears 0% on redmine because a subtask has been added (15442).

Task #15499 It will be finished.

MARGHERITA:

Task #15044: It will be finished; 80% done, now working with Marc for testing.

Task #15343, Task #15346 move to 2.1.0 (they are subtasks of 15058).

STEVE:

Task #15342 move to 2.1.0

Task #15577 It will be finished.

VLADIMIR:

Task #15007 Made code refactoring, moving original code in a function to avoid repetitions.
The errors will be handled now in the function. It should be finished.

Task #15348 move to 2.1.0

Task #15524 It's done but it will be effective in production.

2. Opened tickets

Incident tickets still opened must be updated. Some are finished and should be closed.

Followed discussion about rotation calendar to handle request and incident
Steve: should we go through FIFE support?
Anna: it is my understanding that we should have our rotation calendar (Scientific Production Processing), but at the end we are almost the same people as Distributing Computing Support, Data Handling, etc .. This rise the potential problem having people always on shift.

3. Alex’s requests

Felipe discussed about NOvA request on the time window for active and no-active campaigns (task ...)

4. AOB

Anna: Focus on this week will be the hardware, planning to request new machines.

Steve: It will be important to log more information on the queries executed:
this will help understanding problems in view of optimizing.
The log will grow a lot but we will be able to see how much space is required.

ANNA:
Need to meet with the experiments to help with their workflow and make sure POMS provides what they need for their jobs submission.

G-2: they need some help with managing the proxy so that they don't run into
problems again . Joe will check if their request of managed proxy has been done.
MicroBoone: They are ready , need to meet to make sure all they require is available.
ProtoDune: They are interested in using POMS and CI.
Anna will help to get started with their MC jobs.